OpenStudy (anonymous):

Katie and Pamela put some money into their money boxes every week. The amounts of money (y), in dollars, in their money boxes after certain amounts of time (x), in weeks, are shown by the equations below: Katie y = 30x + 20 Pamela y = 20x + 90 After how many weeks will Katie and Pamela have the same amount of money in their money boxes, and what will that amount be? A. 6 weeks, $150 B. 6 weeks, $230 C. 7 weeks, $150 D. 7 weeks, $230

OpenStudy (welshfella):

this will be when the value of y is the same for both equations so we can say 30x + 20 = 20x + 90 solve for x

OpenStudy (anonymous):

70

OpenStudy (welshfella):

30 - 20x = 90 - 20 10x = 70 x = ?

OpenStudy (anonymous):

7

OpenStudy (welshfella):

right and how can you work out the amount of money after the 7 weeks

OpenStudy (welshfella):

remember y = 2x + 90 is the amount of money saved in x weels so for 7 weeks y = ?

OpenStudy (anonymous):

how do i do the money part?

OpenStudy (welshfella):

just replace the x in 20x + 90 and work it out 20*7 + 90 = ?

OpenStudy (welshfella):

* replace the x by 7

OpenStudy (anonymous):

$230

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

D is the correct option

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

yw